Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $30 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally:

Per Unit

13,000 Units
Per Year

Direct materials

$

13

$

169,000

Direct labor

9

117,000

Variable manufacturing overhead

3

39,000

Fixed manufacturing overhead, traceable

3

*

39,000

Fixed manufacturing overhead, allocated

6

78,000

Total cost

$

34

$

442,000

*One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value).

Required:

1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 13,000 carburetors from the outside supplier?

2. Should the outside supplier’s offer be accepted?

3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $130,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 13,000 carburetors from the outside supplier?

4. Given the new assumption in requirement 3, should the outside supplier’s offer be accepted?
Answer #1

1) Differential analysis

Make Buy
Direct material 169000
Direct labor 117000
variable manufacturing overhead 39000
Fixed manufacturing overhead (39000/3) 13000
Purchase cost (13000*30) 390000
Total relevant cost 338000 390000

Financial (disadvantage) = 338000-390000 = -52000

2) NO

3) Differential analysis

Make Buy
Direct material 169000
Direct labor 117000
variable manufacturing overhead 39000
Fixed manufacturing overhead (39000/3) 13000
Opportunity Cost 130000
Purchase cost (13000*30) 390000
Total relevant cost 468000 390000

Financial advantage = 468000-390000 = 78000

4) Yes